OpenStudy (anonymous):

lim x--> pi/2 of (cosx)/(1-pi)

OpenStudy (anonymous):

= 0 need solution?

OpenStudy (bahrom7893):

Just plug in pi/2 into the expression and evaluate: (Cos(pi/2))/(1-pi) = 0/(1-pi) = 0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):